OpenStudy (anonymous):

\[ Let f:A \rightarrow B $ be a given function. Prove that f is one-to-one (injective) $ \Leftrightarrow f(C\cap D)=f(C)\cap f(D) $ for every pair of sets C and D in A $\]

OpenStudy (anonymous):

Let \[f:A\rightarrow B\] be a given function. Prove that f is one-to-one (injective) \[\leftrightarrow f(C\cap D)=f(C)\cap f(D)\] for every pair of sets C and D in A

OpenStudy (anonymous):

\[Let f:A\rightarrow B be a given function. Prove that f is one-to-one (injective) \\\Leftrightarrow f(C\cap D)=f(C)\cap f(D) for every pair of sets C and D in A\]

OpenStudy (anonymous):

i was just rewriting so i could read it, i am not sure i know how to do it

OpenStudy (anonymous):

well one way is trivial, since \(f(A\cap B)\subset f(A)\cap f(B)\) for any \(f\)

OpenStudy (anonymous):

or does that need clarification as well? we can write it out if you like

OpenStudy (anonymous):

yes we need to right it..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

why letter is not separated

OpenStudy (anonymous):

suppose \(z\in f(A\cap B)\) then \(z=f(x)\) for some \(x\in A\cap B\) making \(x\in A\) and \(x\in B\) so \(z\in f(A)\) and \(z\in f(B)\) therefore \(z\in f(A)\cap f(B)\)

OpenStudy (anonymous):

this shows for any \(f\) you have \(f(A\cap B)\subset f(A)\cap f(B)\)

OpenStudy (anonymous):

now we need to prove that if \(f\) in injective, we have \(f(A\cap B)=f(A)\cap f(B)\) since we already have containment one way, this amounts to showing \[f(A)\cap f(B)\subset f(A\cap B)\]

OpenStudy (anonymous):

pick a \(z\in f(A)\cap f(B)\) so there exists a \(x_1\) in \(A\) with \(f(x_1)=z\) and likewise there is a \(x_2\) in \( B\) with \(z=x_2\) now comes the "injective" part since \(f\) is injective, and \(f(x_1)=f(x_2)=z\) we know \(x_1=x_2\) and so \[z\in f(A\cap B)\]

OpenStudy (anonymous):

typo there, i meant "likewise there exists \(x_2\in B\) with \(f(x_2)=z\) sorry

OpenStudy (anonymous):

so that is the proof one way, that "if \(f\) is injective, then \(f(A\cap B)=f(A)\cap f(B)\)

OpenStudy (anonymous):

other way is easier, since a singleton is a set

OpenStudy (anonymous):

A={x} B={y} you mean like this one..